# A New Approach to the Representation Theory of Lorentzian Pseudo-Tensors

### v1

@@ II.7.2 Vectors
We may for example, denote the momentum representation as $\rho_{+-}$, and this would mean the parity operator acts simply as the defining representations parity operator, while the time reversal operator acts as the defining time reversal operator, with an additional sign flip. Curiously, this table resembles the $\mathcal{P}$ and $\mathcal{T}$ columns of Tab. (1), the list of one dimensional representations. It seems to indicate we might be able to refer to momentum as being in something like the vector version of the $T$ representation, and this would tell us it transforms as in the defining representation, except that operators involving time reversal acquire an additional sign. We will confirm this intuition below.
II.7.3 Tensors
@@
One might have anticipated that higher tensor powers garner us a greater number of possible representations under reflections. This turns out to be false: we will always find exactly four representations of an $n^{\mathrm{th}}$ tensor power of $\operatorname{O}(1,3)$ reps. One can think of this as being due essentially to there only being ‘two’ versions of $\mathcal{P}$ between the four vector representations, and similarly for $\mathcal{T}$. The application of parity inversion on an $n^{\textrm{th}}$ order tensor will apply $n$ parity transformation matrices (of unspecified representation) contracting on the tensors $n$ indices. But these $n$ parity matrices can always be written as $n$ defining parity transformation matrices, and a possibly left over sign. The same argument goes for $\mathcal{T}$. As such the parity inversion or time reversal of any $n^{\textrm{th}}$ order tensor can simply be understood as the application of $n$ defining transforms, and the possible application of a sign.
Taking a look at Tab. (3), hidden in plain sight is the multiplication table of $\operatorname{K}_{4}$, the Klein four group, our discrete reflection group! It becomes explicit if we perform the identification of names of representations given in Tab. (4).
@@
Table 4: Generic $\operatorname{O}(1,3)$ tensors also show up in ‘reflection quartets’ which can be labeled just as the one dimensional representations. The details of the reflections are still more complicated than the one dimensional case, but the tensor products will follow the familiar pattern of $\operatorname{K}_{4}$.
What we see from this is that a tensor product of a coordinate type vector (‘$1$ type vector’) with a momentum type vector (‘$T$ type vector’), will give a ‘$T$ type second order tensor’, i.e. this type of tensor transforms as the defining rep, but under any time reversal the components acquire an additional overall sign. This enormously simplifies keeping track of the transformation properties of tensors components: it is in principle no more complicated than the $1$ dimensional representations. Note also this assignment of names agrees with the scalar case: $x^{\mu}p_{\mu}$ is a scalar which is odd under $\mathcal{T}$, and so it takes the scalar representation $T$. Let it be made explicit however that when saying a tensor or vector takes some representation ‘$T$’ or ‘$PT$’ etc, we are not merely saying that the whole tensor picks up a sign under those transforms. We are saying that when ‘naively’ applying the defining parity or time reversal matrices to the object, those particular transforms must pick up an additional sign in order actually be the correct transform of that object. Mathematically of course this is the same as treating the representation theory as if there are only $\mathcal{P}_{\mathrm{def}}$ and $\mathcal{T}_{\mathrm{def}}$, and as if the (for e.g.) $T$ charge simply acts as in the one dimensional case picking up a sign under time reversals. The appeal of this mode of thinking is that we may largely forget about all the ways $\mathcal{P}$ and $\mathcal{T}$ can act on the representations — we need only remember the defining reps, and the $1$-dimensional representations of the Klein group. However it is important to keep in mind that the seemingly ‘extra’ signs are absolutely necessary to correct for using the operators from the wrong representation; the apparently ‘extra’ signs do not enter via an ad hoc extension.
II.8 Examples
II.8.1 Raising and Lowering indices
It is worth seeing ‘why’ for example, $1$-type tensors deserve the name. The metric tensor is a $1$-type tensor: this is reassuring since we do not anticipate that the parity and time reversal properties should be affected by the raising and lowering of indices. We can see this explicitly with a $T$ type tensor (such as the Faraday or angular momentum tensors). One can readily verify, in terms of the one dimensional representations of the components:
@@ II.8.1 Raising and Lowering indices
This is an explicit verification of the general rule for all tensors of various type that $1\cdot T\cdot 1=T$. Interaction with $1$-type tensors in general will leave us in the same type of reflection representation, regardless of the underlying $\operatorname{SO}^{+}(1,3)$ structure.
II.8.2 An Observers $\mathbf{E}$ Field
@@
Without digging into the components, we already know the resulting vector is a $1$-type vector, a.k.a a coordinate type vector, and so its parity and time reversal properties are transparent to us without the need to pick apart its various pieces. For reference, in the rest frame this simply gives: $E^{\mu}=(0\>\>\mathbf{E})^{T}$, which being a coordinate type vector solidifies our understanding that electric fields are odd under parity and even under time reversal.
II.8.3 Pauli–Lubanski Pseudo-Vector & The Levi-Civita Symbol
Typically this object is most useful when $M$ and $p$ are taken to be operators, and values of the scalar $W^{\mu}W_{\mu}$ are used to classify different representations of the Poincare group Weinberg_1995 . Here we have two objects for which we do not know their transformation properties under the reflection group, $W$ and $\varepsilon$. However, we understand the transformation properties of $M$ and $p$ and so after computation can figure out by inspection the transformation properties of the components of $W$, which can then be used to tell us about the Levi-Civita symbol. The first component, $W_{0}$, is odd under parity inversion, and overall even under time reversal, as both $\mathbf{J}$ and $\mathbf{p}$ are $\mathcal{T}$-odd. This means $W_{0}$ is charged under the reflection group as $P$ — by our classification of vectors, this means it had better be the case the spacial part of $W$ is charged as $T$. Inspection shows that is true. So we have confirmed that $W^{\mu}$ is an axial type vector, i.e a $PT$ type vector, and will ‘pick up a sign’ under each of time and space reflections.
The Levi-Civita symbol transforms just as any four index tensor, however it picks up signs under space or time reflections. This is not the fact it appears to be on the surface. As $\varepsilon^{\mu\nu\sigma\rho}$ is the invariant tensor which delivers us determinants, for which the reflections $\mathcal{P}_{\mathrm{def}}$ and $\mathcal{T}_{\mathrm{def}}$ have $-1$, the charge under the reflection group will cancel these signs out. Thus it is the case that in fact $\varepsilon$ is exactly invariant under the entire Lorentz group. [^10] It is worth mentioning this fact about $\varepsilon$ is typically understood via the concept of tensor densities (see Weinberg_1972 ), however the purely representation theoretic perspective here gives a much more satisfying understanding of the Levi-Civita symbol’s properties.
II.8.4 An Observers $\mathbf{B}$ Field
@@
II.9 Maxwell’s Equations are ‘Reflection Homogeneous’
@@ II.9.1 Parity Reflected Solutions
II.9.2 Time Reflected Solutions
@@
II.10 The Heat Equation is not ‘Reflection Homogeneous’
@@
III Summary
A modern representation theory of the entire Lorentz group $\operatorname{O}(1,3)$ has been presented. In particular, it was shown that $\operatorname{O}(1,3)=\operatorname{SO}^{+}(1,3)\rtimes\operatorname{K}_{4}$. The nature of this semi-direct product constrained the kinds of four vector representations allowed, and in turn we were able to argue that there can be precisely $4$ different transformation properties for any $n^{th}$ order tensor power of the entire Lorentz group. The four distinct representations of any tensor power can be labeled $(1,P,T,PT)$. The label of the representation tells us precisely how the tensor’s transformation under reflections, differs from the defining representation. A product of two objects will be in the representation of $\operatorname{K}_{4}$ labeled by the result of the multiplication of the labels under the Klein group. So for example the representation of the Faraday tensor $F^{\mu\nu}=x^{\mu}p^{\nu}-x^{\nu}p^{\mu}$ in the group $\operatorname{O}(1,3)$ would be computed as
As representations of $\operatorname{SO}^{+}(1,3)$ can be labelled by pairs of half integers, representations of $\operatorname{O}(1,3)$ can be labeled by pairs of half integers, alongside a charge under the discrete reflection group $\operatorname{K}_{4}$, which is interpreted as above. Finally several examples were constructed to show off the utility of the formalism in reading off the charges under the reflection group of different objects, so as the make the behavior of any tensorial object under arbitrary reflections transparent to us. It was also shown that equations which are homogeneous in their $\operatorname{K}_{4}$ charges exhibit parity and time reversal symmetry, whereas those which are not cannot exhibit these symmetries.
IV Author Declarations
The author has no conflicts to disclose.
Appendix A The Semi-Direct Product Structure
@@
This looks mostly like the law for a direct product of group actions, except that the group multiplication for the $G$ component has been ‘twisted’ by $\psi$. This may be interpreted as a commutation relation: the effect of moving $h_{2}$ ‘past’ $g_{1}$ results in $g_{1}$ being altered in the way defined by $\psi$. It was seen above that this behaviour is precisely how the discrete reflection group $\operatorname{K}_{4}$ interacts with the proper orthochronus Lorentz group.
Appendix B Automorphisms of the Lorentz Group

### v2

@@ II.6.1 Scalars and Vectors
We may for example, denote the momentum representation as $\rho_{+-}$, and this would mean the parity operator acts simply as the defining representations parity operator, while the time reversal operator acts as the defining time reversal operator, with an additional sign flip, the intention being to forget about all the various incarnations of the operators, and only have to remember the defining ones, and then keep track of some signs.
Curiously Tab. (2) resembles the $\mathcal{P}$ and $\mathcal{T}$ columns of Tab. (1), the list of one dimensional representations of $\operatorname{K}_{4}$. In fact we could just as well label our scalar representations by their $\mathcal{P}$ and $\mathcal{T}$ parities just as well, identical to this renaming of the vector case. This seems to indicate we might be able to refer to momentum as being in something like the vector version of the $T$ representation, and this would tell us it transforms as in the defining representation, except that operators involving time reversal acquire an additional sign. We will confirm this below, that there are four representations of $\operatorname{O}(1,3)$ for every representation of $\operatorname{SO}^{+}(1,3)$, and these representations may be labeled by elements of the Klein group which in turn indicate additional sign changes garnered under the basic spacetime reflections.
II.6.2 Tensors
@@
One may have anticipated that higher tensor powers, by their having a greater number of components, would garner us a greater number of possible representations under reflections; this is false. We will always find exactly four representations of an $n^{\mathrm{th}}$ tensor power of $\operatorname{O}(1,3)$ reps, distinct in their spacetime reflection properties. The application of parity inversion on an $n^{\textrm{th}}$ order tensor will apply $n$ parity transformation matrices (of unspecified representation) contracting on the tensors $n$ indices. But these $n$ parity matrices can always be written as $n$ defining parity transformation matrices, and a possibly left over sign. Turning this around we may say parity inversion of any $n^{\textrm{th}}$ order tensor is the application of $n$ defining transforms, and the possible application of a sign. An identical argument holds for $\mathcal{T}$.
Taking a look at Tab. (3), hidden in plain sight is the multiplication table of $\operatorname{K}_{4}$, our discrete reflection group! It becomes explicit if we perform the identification of names of representations given in Tab. (4).
@@
Table 4: Generic $\operatorname{O}(1,3)$ tensors show up in ‘reflection quartets’ which can be labeled just as the one dimensional representations. The details of the reflections are still more complicated than the one dimensional case, but the tensor products will follow the same pattern of $\operatorname{K}_{4}$.
It is easy to see the right side version of Tab. (3) remains true no matter the underlying representation of $\operatorname{SO}^{+}(1,3)$, and so every representation of $\operatorname{SO}^{+}(1,3)$ will have four corresponding representations in $\operatorname{O}(1,3)$, one of each type $1,P,T,PT$. We see from this, for example, that a tensor product of a coordinate type vector (‘$1$ type vector’) with a momentum type vector (‘$T$ type vector’), will give a ‘$T$ type second order tensor’, i.e. this type of tensor transforms as the defining rep, but under any time reversal the components acquire an additional overall sign. This enormously simplifies keeping track of the transformation properties of tensors components: it is in principle no more complicated than the $1$ dimensional representations. This assignment of names agrees with the scalar case: $x^{\mu}p_{\mu}$ is a scalar which is odd under $\mathcal{T}$, and it takes the scalar representation $1\cdot T=T$.
It bears repeating that when saying a tensor or vector takes a representation ‘$T$’ or ‘$PT$’ etc, we are not merely saying that the whole tensor picks up a sign under those transforms. We are saying that when ‘naively’ applying the defining parity or time reversal matrices to the object, those particular transforms must pick up an additional sign in order actually be the correct transform of that object. Mathematically of course this is the same as treating the representation theory as if there are only $\mathcal{P}_{\mathrm{def}}$ and $\mathcal{T}_{\mathrm{def}}$, and as if the (for e.g.) $T$ charge simply acts as in the one dimensional case picking up a sign under time reversals. The appeal of this mode of thinking is that we may largely forget about all the ways $\mathcal{P}$ and $\mathcal{T}$ can act on the representations — we need only remember the defining reps, and the $1$-dimensional representations of the Klein group. However it is important to keep in mind that the seemingly ‘extra’ signs are absolutely necessary to correct for using the operators from the wrong representation; the apparently ‘extra’ signs do not enter via an ad hoc extension.
II.7 Important Examples
II.7.1 The Metric: Raising and Lowering Indices
It is worth seeing why $1$-type tensors deserve the name. The metric tensor is a $1$-type tensor: this is reassuring since we do not anticipate that the parity and time reversal properties should be affected by the raising and lowering of indices. We can see this for example with a $T$ type tensor, such as the Faraday or angular momentum tensors. One can readily verify, in terms of the one dimensional representations of the components:
@@ II.7.1 The Metric: Raising and Lowering Indices
This is an example of how the multiplication of the Klein group representations (here $1\cdot T\cdot 1=T$) may be carried out no matter the underlying $\operatorname{SO}^{+}(1,3)$ structure of the tensors involved. Interaction with $1$-type tensors in general will leave us in the same type of reflection representation.
II.7.2 The Levi-Civita Symbol & Pauli–Lubanski Pseudo-Vector
Typically this object is most useful when $M$ and $p$ are taken to be operators, and values of the scalar $W^{\mu}W_{\mu}$ are used to classify different representations of the Poincare group Weinberg_1995 . Here we have two objects for which we do not know their transformation properties under the reflection group, $W$ and $\varepsilon$. However, we understand the transformation properties of $M$ and $p$ and so after computation can figure out by inspection the transformation properties of the components of $W$, which can then be used to tell us about the Levi-Civita symbol. The first component, $W_{0}$, is odd under parity inversion, and overall even under time reversal, as both $\mathbf{J}$ and $\mathbf{p}$ are $\mathcal{T}$-odd. This means $W_{0}$ is charged under the reflection group as $P$ — by our classification of vectors, we expect the spacial part of $W$ to be charged as $T$. Inspection shows that is true. So we have confirmed that $W^{\mu}$ is an axial type vector, i.e a $PT$ type vector, and will ‘pick up a sign’ under each of time and space reflections.
The Levi-Civita symbol transforms just as any four index tensor, however it picks up signs under space or time reflections. This may not be the fact it appears to you as on the surface: as $\varepsilon^{\mu\nu\sigma\rho}$ is the invariant tensor which delivers us determinants, for which the reflections $\mathcal{P}_{\mathrm{def}}$ and $\mathcal{T}_{\mathrm{def}}$ have $-1$, the charge under the reflection group will cancel these signs out. Thus it is the case that in fact $\varepsilon$ is exactly invariant under the entire Lorentz group. [^8] It is worth mentioning this fact about $\varepsilon$ is typically understood via the concept of tensor densities (see Weinberg_1972 ). I believe this purely representation theoretic perspective gives a much more satisfying understanding of the Levi-Civita symbol’s properties.
Summary and Conclusions
A convenient representation theory of the entire Lorentz group $\operatorname{O}(1,3)$ has been constructed. In particular the decomposition $\operatorname{O}(1,3)=\operatorname{SO}^{+}(1,3)\rtimes\operatorname{K}_{4}$, understood as the fact that boost parameters are charged under parity and time reversal, constrains the representations of the discrete reflection group $\operatorname{K}_{4}$ upon four-vectors to be one of four combinations, and subsequently it was shown there are precisely four distinct representations of $\operatorname{O}(1,3)$ for each representation of $\operatorname{SO}^{+}(1,3)$. These four distinct representations can be labeled $1,P,T,$ and $PT$; and the label of the representation tells us precisely the tensor’s behavior under spacetime reflections $\mathcal{P}$ and $\mathcal{T}$. Specifically, all tensors may be transformed by the defining (coordinate) representation of the parity and time reversal operators, and those in the $P,T$, or $PT$ representations gain an additional sign under the transforms in their names.
The representation of a product of any two tensors, whether through an inner product, tensor product, contraction etc, will be labeled by the multiplication of the $\operatorname{K}_{4}$ labels of each factor, completely agnostic to the particular $\operatorname{SO}^{+}(1,3)$ structure of the tensors at play. So for example the contraction of a $PT$ type 2-tensor with a $T$ type vector results in a $P$ type vector. In toto, in this formalism the representation of say, the Faraday tensor $F^{\mu\nu}=x^{\mu}p^{\nu}-x^{\nu}p^{\mu}$ in the full group $\operatorname{O}(1,3)$ would be given as
Since representations of $\operatorname{SO}^{+}(1,3)$ can be labeled by pairs of half integers, representations of $\operatorname{O}(1,3)$ can be labeled by pairs of half integers, alongside a charge under the discrete reflection group $\operatorname{K}_{4}$, interpreted as above. Several further examples are given Appendix. (A) to show off the utility of the formalism. By keeping track of a simple label the behavior of any tensorial object under arbitrary reflections is made transparent to us.
Future Work
The author intends to pursue a similar project for the spinoral double coverings of $\operatorname{O}(1,3)$, the $\operatorname{Pin}$ groups, in order to further study, classify, and clarify the relationship between Lorentz invariance, parity violation, and the $\mathcal{CPT}$ theorem.
Author Declarations
The author has no conflicts to disclose. Data sharing is not applicable to this article as no new data were created or analyzed in this study.
Appendix A Further Examples
A.1 The Faraday Tensor and the Potential $A^{\mu}$
While the four potential $A^{\mu}$ is not a 2-tensor, a discussion of the Faraday tensor is incomplete without first addressing it. The parity inversion properties of $A^{\mu}$ are rarely up for debate. It is the time reversal of $A^{\mu}$ which is the cause of contention in the literature. In particular, it is the question of whether a time reversal should ‘include’ a charge conjugation. This debate is due in part to the Feynman–Stückelberg interpretation of anti-particle solutions, and in part due to the anti-unitarity of the time reversal operator. However, so long as we are considering the case of classical fields, we can cleanly separate time reversal from charge conjugation without debate. In what follows, the time reversal of a stationary charge $q$ will be understood to produce a description of an identical system, and not a description of a system of a stationary charge of $-q$. This can be stated as the postulate that classical electrical charges live in the trivial representation of the discrete reflection group.
It is not clear a priori which vector rep of $\operatorname{O}(1,3)$ the potential $A^{\mu}$ takes. One could argue from its equations of motion that it should have the same properties as a four-current; but this requires first assuming parity and time reversal invariance of those equations. The answer can be gleaned from the much less assumptive but important expectation that the canonical momentum of the electromagnetic field has well defined transformation properties under the discrete reflection group for all observers. The canonical momentum which enters the Hamiltonian of the electromagnetic field is given by JACKSON3RD :
$P^{\mu}=\left(p^{\mu}+eA^{\mu}\right),$ (35)
where $p^{\mu}$ is the mechanical momentum of the relevant particle with charge $e$. If observers related by (proper orthochronus) Lorentz transforms are to agree about the charges of $P^{\mu}$ under the discrete reflection group, it must be the case that $K[eA^{\mu}]=K[p^{\mu}]=(m)$. Committing to our assumption that $e$ is invariant under the reflection group, we must have $K[A^{\mu}]=K[p^{\mu}]=1\oplus\vec{PT}=(m)$, i.e. $A^{\mu}$ is a momentum type vector.
This representation of $\operatorname{O}(1,3)$ is identical to that of the angular momentum tensor:
$K[F^{\mu\nu}]=\begin{pmatrix}T\\ \vec{P}\end{pmatrix}\otimes\begin{pmatrix}1\\ \vec{PT}\end{pmatrix}=\begin{pmatrix}T&\vec{P}\\ \vec{P}&T\mathbb{I}_{3}\end{pmatrix}.$ (37)
As such, parity and time reversal will act the same here as they did upon the angular momentum tensor:
This is what is taught in most physics textbooks, and finally we can see a consistent representation theoretic defense of it. The only room for disagreement is the debate of whether or not charges should take non-trivial representations of the discrete reflection group, which presently seems a discussion better suited for realm of quantum field theory. We may also see clearly why what Arntzenius & Greaves GREAVESARNTZENIUS dubbed ‘Malaments proposal’ works:
It is clear why the time reversal of the Faraday tensor should look like the application of parity, followed by an overall additional minus sign on the whole tensor.
A.2 An Observers $\mathbf{E}$ Field
@@
Without digging into the components, we already know the resulting vector is a $1$-type vector, a.k.a. a coordinate type vector, and so its parity and time reversal properties are transparent to us without the need to pick apart its various pieces. For reference, in the rest frame this simply gives: $E^{\mu}=(0\>\>\mathbf{E})^{T}$, which being a coordinate type vector solidifies our understanding that electric fields are odd under parity and even under time reversal.
A.3 An Observers $\mathbf{B}$ Field
@@
A.4 Maxwell’s Equations
@@ A.4.1 Parity Symmetry
A.4.2 Time Reversal Symmetry
@@
A.5 The Heat Equation
@@
Appendix B $\operatorname{O}(1,3)$ as a Semi-direct Product Group
B.1 Semi-Direct Products
@@
This looks mostly like the law for a direct product of group actions, except that the group multiplication for the $G$ component has been ‘twisted’ by $\psi$. This may be interpreted as a commutation relation: the effect of moving $h_{2}$ ‘past’ $g_{1}$ results in $g_{1}$ being altered in the way defined by $\psi$. It was seen above that this behavior is precisely how the discrete reflection group $\operatorname{K}_{4}$ interacts with the proper orthochronus Lorentz group.
B.2 Outer Automorphisms of the Lorentz Group
@@
B.3 Semi-Direct Product Structure of $\operatorname{O}(1,3)$
On a generic Lorentz transform, $\mathcal{P}$ and $\mathcal{T}$ leave the rotation angles the same, but reverse the boost parameters. This action is equivalent to saying we have acted via an outer automorphism $\mathcal{\hat{O}}$ upon our Lie group: [^14] See appendix (B.2) for review of the outer automorphism group of the proper orthochronus Lorentz group
This is precisely the behavior of a semi-direct product (see appendix (B.1) for review): commuting a first kind of group element past a second kind, results in an automorphism being applied to the first kind of group element. This observation tells us explicitly how to build the semi-direct product. We are seeking a map $\psi$ that takes elements of the discrete reflection group, and maps them to automorphisms of $\operatorname{SO}^{+}(1,3)$:
Let $\psi$ be the following map: